Find mean & margin of error if 28.4 < µ < 38.8


1
Expert's answer
2021-05-07T09:48:54-0400

Mean "= \frac{28.4+38.8}{2}=33.6"

Margin of error "= \frac{38.8 -28.4}{2} = 5.2"